### v1

@@
A direct method for reproducing fully relativistic spectra from standard accretion disks by modifying their inner boundary
@@
Aims. A proper relativistic treatment would be troublesome, unappealing, or require more advanced computational tools (e.g. the kerrbb code). Our goal is to fully incorporate the black hole’s spin and all the general relativistic effects on the observed spectra coming from X-ray binary systems while maintaining the simplicity of the standard disk model. That is possible by appropriately shifting only the disk’s inner radius.
Methods. We employ some of the most efficient pseudo-Newtonian potentials around Kerr black holes and derive two generalized disk temperature profiles, thus incorporating the spin’s contribution to the thermal spectra. Then, we associate the observed radiative efficiency with the emission pattern featuring all the relativistic effects included in the kerrbb model, obtaining an expression about the modified inner radius of the disk. Moreover, we apply this method to Cygnus X-1 by fitting the observational data obtained during its high/soft and hard/low spectral states.
Results. The fully relativistic spectra are reproduced to a very good approximation with an error margin of 0.03-4% by a standard disk model with a disk temperature expression emerging from a Kerr pseudo-Newtonian potential. The disk is parameterized by a modified innermost radius within the range of $(0.2-2)R_{ISCO}$, depending on the source’s viewing angle and black hole spin. That treatment compensates for the error introduced through the potential and incorporates the correct angular dependence of the spectra. Relativistic effects near the black hole make an otherwise standard accretion disk with inclination $\theta<60^{\circ}$ seem truncated to larger radii to a distant observer. On the other hand, an edge-on view of the disk gives the perspective of being pulled closer to the central object than the respective ISCO radius. In addition, we show that the observational data of Cygnus X-1 can be satisfactorily fitted by employing a reasonably simple lepto-hadronic jet model and a hybrid thermal/non-thermal corona along with the Kerr-adjusted standard accretion disk.
@@
XRB - accretion disk - spin - pseudo-Newtonian - general relativity - ISCO - Cygnus X-1
@@ 1 Introduction
The numerical code kerrbb integrates all the relativistic effects, including the disk’s self-irradiation caused by light deflection (Li et al., 2005). It is employed as part of the xspec package (Arnaud, 1996) for fitting accretion disk spectra through a ray-tracing method providing a black-body emission continuum. Nonetheless, finding a straightforward and simpler way to estimate the fully relativistic spectra from a standard accretion disk is necessary for constructing relativistic-equivalent models to the Shakura-Suniaev prototype.
In this work, we employ two efficient pseudo-Newtonian (PN) gravitational potentials that approximate the boundary conditions at the edge of the accretion disk around a rotating black hole. We aim to develop valid temperature profiles that outline the thermal energy distribution across the accretion disk surface. Furthermore, we replicate the correct spectral angular distribution by appropriately adjusting the disk’s innermost boundary. The latter is achieved based on the emission pattern, as determined by the observed radiative efficiency of Kerr black holes following the kerrbb model (Campitiello et al., 2018). Moreover, this approach enables us to avoid the complexities of a complete general relativistic treatment akin to that of Novikov & Thorne (1973). Additionally, it solidifies a method for fully integrating general relativity in fitting the observed spectra from XRBs while upholding the fundamental assumptions and simplicity of the Shakura-Suniaev model.
@@
Cygnus X-1 is one of the most well-studied and observed XRBs in our Galaxy. Measurements of its spectral distribution in the soft state by RXTE, ASCA, and OSSE in 1996 (Ling et al., 1997; Dotani et al., 1997) indicate a thermal component from the accretion disk and a hard tail up to the MeV scale. The latter is assumed to originate from the up-scattering of seed photons from the disk in a thermal/non-thermal corona hybrid.
@@
In the remainder of this work, we discuss the disk geometry and the radial temperature obtained from the two PN solutions referring to rotating black holes in Sec. 2. In Sec. 3, we discuss the radiative efficiency perceived by an observer and how it can be employed to acquire an expression for the location of the disk’s inner edge as a function of the spin and the viewing angle. Then, in Sec. 4, we apply this method to the multi-wavelength spectra detected from Cygnus X-1 in its two primary spectral states. Finally, we conclude our results in Sec. 5.
@@ 2.2 The pseudo-Newtonian solution to the standard accretion disk
where $r=R/R_{g}$, $r_{h}=1+\sqrt{1-\alpha_{*}^{2}}$ is the event horizon in the Kerr metric, and $\beta=r_{in}/r_{h}-1$ denotes a dimensionless constant for a given specific black hole angular momentum. The respective integrated potential reproduces pretty accurately the location of the marginally stable orbit and the disk radiative efficiency as predicted by general relativity.
@@ 2.3 Spin-dependent disk temperature profiles
In comparison, Eqs. (6) produces negligibly larger disk luminosity across the spectrum for energies below 1 keV that drops slightly more abruptly for higher energies. The deviation is marginal for large $\alpha_{*}$ in the co-rotation case, while they coincide for the rest of the spin values. Both expressions reduce to the temperature function of Gierliński et al. (1999) that implements the Paczyński-Wiita potential.
@@
Extreme Kerr black holes (i.e., $\alpha_{*}>0.9$) are verified as the compact objects for an increasing number of well-known XRBs (McClintock et al., 2006; Stuchlík & Kološ, 2016; Zhao et al., 2021). The frame-dragging effects grow much more important, approaching the event horizon. Therefore, the spin-dependent profiles outlined in Eqs. (6) and (9) yield a more comprehensive understanding of the physical processes governing the accretion flow around a black hole’s immediate vicinity.
@@ 3 Spectral angular distribution via disk surface fluctuations
The detectable spectra from the accretion disk are heavily dependent on the observer’s viewing angle. Ignoring all relativistic effects, the observed radiative efficiency is $\eta_{rad}^{obs}=2cos\theta\eta_{rad}(R_{ISCO})$, where $\eta_{rad}=L_{disk}/\dot{M}c^{2}$. Regardless, frame-dragging, gravitational redshift, Doppler boost, and light-bending effects significantly alter the angular dependence of the perceived emission flows near the black hole. Hence, the emerging emission pattern following the relativistic kerrbb model is determined by an analytic expression derived by Campitiello et al. (2018) as $\eta_{rad}^{obs}=f(\theta,\alpha_{*})\eta_{rad}(R_{ISCO})$. It is given in the Appendix B and is thoroughly demonstrated for varying values of $\alpha_{*}$ in Fig. 4. We also include the radiative efficiency produced with both pseudo-potentials discussed in this work and the respective theoretical predictions of Novikov-Thorne. The comparison between the PN adaptation and general relativity reveals an increasing deviation for highly rotating black holes.
@@
Figure 4: The solid lines correspond to the radiative efficiency given by the analytic counterpart of the kerrbb model as observed for different system inclinations and black hole spin values (referring to Campitiello et al. (2018) as C18). The purple dashed line showcases the angular dependence corresponding to a Shakura-Suniaev disk around a static black hole, neglecting all relativistic contributions. The colored dotted lines for each spin value depict the respective theoretical predictions of Novikov & Thorne (1973). For comparison, the respective black and orange-dotted lines represent the Kerr pseudo-Newtonian treatment of this paper.
@@ 3.1 Modification of the disk’s inner boundary
The radiative efficiency is independent of the accretion rate and the spectral hardening factor $f_{col}$. It depends negligibly on the black hole’s stellar mass $M_{bh}$ and the disk’s inner radius $R_{in}$, as long as the outer radius is large enough, $R_{out}>10^{11}\ cm$, akin to the default value set in kerrbb. Fig. 5 illustrates this correlation, which is established by employing Eqs. (6) and (9) in calculating the thermal spectra from the disk. In this work, we fit the power-law form (black dotted lines in Fig. 5), particularly distinguishable in the case of the Artemova solution, with the following expression
@@
The respective orange solid lines are fitted using a power-law index of $p=0.75$ instead. A noticeable deviation for all spin values, especially near the ISCO radius, deter us from further exploiting the Mukhopadhyay pseudo-force. Also, the fact that the Artemova potential works for radii significantly smaller than $R_{ISCO}$ makes it a very suitable tool for matching the fully relativistic spectrum for a given angle of view.
@@
Furthermore, incorporating the correct angular distribution urges the equivalence between Eq. (14) and the observed efficiency given by the analytic function of Campitiello et al. (2018) (see Appendix B). By doing so, we find an expression for the modified inner boundary of the accretion disk, which is given by
@@
This way, the fully relativistic spectra are reproduced by a standard disk model with the temperature profile of Eq. (6) and an innermost radius given by Eq. (15). The latter is plotted in Fig. 6 and depends on the inclination and angular momentum of the accreting compact object. Fig. 6 shows that, independently of the rotation, the physical processes associated with an accretion disk observed from an angle of approximately $\theta=60^{\circ}-65^{\circ}$ identify with the Newtonian-like dynamics of the Shakura-Suniaev model. Thus, a minimum shift of $R_{in}$ is required to fit the corresponding kerrbb spectra. Consequently, the boundary-modification approach is exceptionally accurate regarding this angle range.
@@
Figure 6: Modified innermost radii integrating the disk’s fully relativistic treatment of the kerrbb model, which accounts for the viewing angle and the black hole’s rotation.
In Fig. 7, we present the energy distributions produced in the regions near a static and rapidly rotating black hole, adopting the boundary-shifting approach. Those results are in a very good agreement with the fully relativistic formalism of Novikov and Thorne incorporated in the kerrbb model, as seen in Fig. 3 of Campitiello et al. (2018) and Reynolds (2021) (Fig. 4, p. 131).
@@
Figure 7: Fully relativistic spectra emerging from adjusting the inner disk radius for the two extreme cases of $\alpha_{*}=0$ and $\alpha_{*}=0.998$ and various inclinations of the disk’s plane.
4 Results and discussion
@@
References. Miller-Jones et al. (2021), Brocksopp et al. (1999), Krawczynski et al. (2022), Stirling et al. (2001), Gou et al. (2014), Tetarenko et al. (2019)
4.1 High/Soft spectral state
Using Eqs. (6), (15), and the recently redefined Cygnus X-1 parameters of Table 1, we showcase the spin-dependent disk spectra while the system is going through the soft state in Fig. 8. The solid lines correspond to the alignment between the black hole and accretion disk rotational directions. It should be mentioned that those results were obtained for a fixed mass accretion rate of $\dot{M}=8.5\times 10^{-8}$ $M_{\odot}$/yr, which, in most cases, is a free model parameter. In addition, we estimate the radiative efficiency of an extreme rotating black hole to be $\eta_{rad}\approx 0.35$, which agrees with the theoretical predictions. Moreover, employing this work’s pseudo-Newtonian and boundary-shifting approaches allows us to constrain more accurately the black hole’s accretion rate as long as we have a precise value of the spin or vice versa.
@@ 4.1 High/Soft spectral state
In Fig. 9, we outline the data obtained during observations of the soft spectral state of Cygnus X-1 (Ling et al., 1997; Dotani et al., 1997). The thermal emission at lower energies is fitted by a black-body spectrum emitted by the inner disk regions. The respective spin is taken as $\alpha_{*}=0.992$, satisfying the lower limit set by Miller-Jones et al. (2021) for the slightly misaligned spin-orbit case. That corresponds to an accretion disk close to the compact object (i.e., $R_{in}\approx R_{g}$), a primary characteristic of the high/soft state.
@@
$\dot{m}$ 0.006 0.004 mass accretion rate
@@
$T_{cor}$ [keV] 60 150 corona temperature
$\tau_{cor}$ 0.26 0.50 thermal corona optical depth
p 4.3 1.8 corona power-law index
$\gamma_{min}$ 2.2 1.7 minimum energy of non-thermal electrons
@@
h 0.086 0.294 non-thermal/thermal electrons ratio in the corona
@@
$n_{p}/n_{e}$ 0.009 ratio of hadrons to leptons
@@
Moreover, the hard tail above 5 keV is reproduced by the inverse-Compton (IC) scattered disk emission off a thermal/non-thermal corona hybrid described by the parameters of Table 2. In this case, the corona is much more weakly present compared to the hard/low state. The power-law spectrum indicates an electron (or positron) distribution injection, differentiating the two leading emission mechanisms near 10 keV. The injected distribution could be a product of an intense photon annihilation process above the disk level that softens by gradually losing energy due to IC scattering. The resulting power-law has an index of $p>3$ since $p_{inj}>2$. We neglect Compton reflection of the scattered spectrum and irradiation effects beyond those integrated into the kerrbb model as they overcome the purpose of this work.
We assume the injection of 86 non-thermal electrons blending in a thousand thermal ones. Those particles preserve their initial power-law shape, while the rest are described by a Maxwell–Jüttner distribution, which dominates for energies below $\gamma=2.2$. The optical depth in Table 2 corresponds to the thermal counterpart of the corona hybrid. Assuming a fixed spherical region, injecting power-law particles increases corona’s opacity by a factor of $1+h$, where $h$ is the fraction of accelerated to thermal electrons (or positrons) (i.e., $\tau_{cor}^{h}=\tau_{cor}(1+h)\approx 0.28$ in our case).
@@
Figure 9: The high/soft spectral state of Cygnus X-1 was observed by RXTE, ASCA, and OSSE (Ling et al., 1997; Dotani et al., 1997). The observational data are fitted by a black-body spectrum emitted by the inner disk region and an IC scattered component of the disk emission off a combination of thermal and non-thermal corona electrons.

### v2

@@
A simplified approach for reproducing fully relativistic spectra in X-ray binary systems: Application to Cygnus X-1
@@
Aims. A proper relativistic treatment would be troublesome, unappealing, or require more advanced computational tools (e.g. the kerrbb code). Our goal is to fully incorporate the black hole’s spin and all the general relativistic effects on the observed spectra coming from X-ray binary systems while maintaining the simplicity of the standard Shakura-Sunyaev disk model.
Methods. We propose a way to replicate general relativistic spectra as predicted by the Novikov-Thorne model and the kerrbb numerical code by assuming a standard accretion disk with a shifted inner boundary that depends on the black hole spin and the source’s viewing angle. An essential aspect in employing this approach for a broader range of disk inclinations is the derivation of spin-dependent generalized temperature profiles for the accretion disk, obtained from some of the most efficient pseudo-Newtonian potentials around Kerr black holes. We then apply this method to Cygnus X-1, fitting the observational data obtained during its high/soft and hard/low spectral states.
Results. The fully relativistic spectra are reproduced to an excellent approximation with an error margin of 0.03-4% by a standard disk model with a modified innermost radius within the range of $R_{in}=(0.2-2)R_{ISCO}$, depending on the source’s viewing angle and black hole spin. This approach produces observed spectra as predicted by general relativity without the need for the ray-tracing method and complex numerical calculations. Thus, it emerges as a more straightforward alternative way of estimating black hole spins through continuum-fitting by successfully blending the general relativity properties with the Newtonian simplicity in a more complete way than the pseudo-Newtonian solutions. Relativistic effects near the black hole make an otherwise standard accretion disk with inclination $\theta<60^{\circ}$ seem truncated to larger radii to a distant observer. On the other hand, an edge-on view of the disk gives the perspective of being pulled closer to the central object than the respective ISCO radius. In addition, we show that the observational data of Cygnus X-1 can be satisfactorily fitted by employing a reasonably simple lepto-hadronic jet model and a hybrid thermal/non-thermal corona along with the relativistic-equivalent standard thin accretion disk.
@@
XRB - accretion disk - general relativity - spin - pseudo-Newtonian - ISCO - Cygnus X-1
@@ 1 Introduction
The numerical code kerrbb integrates all the relativistic effects according to the Novikov-Thorne thin disk model (Novikov & Thorne, 1973), including the disk’s self-irradiation caused by light deflection (Li et al., 2005). It is employed as part of the xspec package (Arnaud, 1996) for fitting accretion disk spectra through a ray-tracing method providing a black-body emission continuum. Nonetheless, finding a straightforward and simpler way to estimate the fully relativistic spectra from a standard accretion disk is necessary for constructing relativistic-equivalent models to the Shakura-Sunyaev prototype.
In this work, we aim to reproduce the fully relativistic disk spectra and their dependence on the black hole spin and disk inclination by appropriately modifying the disk’s innermost boundary. The latter is achieved based on the emission pattern, as determined by the observed radiative efficiency of Kerr black holes following the kerrbb model (Campitiello et al., 2018). This approach requires implementing efficient pseudo-Newtonian (PN) gravitational potentials that are applicable even at smaller radii than the Innermost Stable Circular Orbit (i.e., ISCO) limit. Hence, we need to derive valid temperature profiles that depend on the black hole’s angular momentum.
Moreover, this process enables us to avoid the complexities of a complete general relativistic treatment akin to that of Novikov & Thorne (1973). Additionally, it solidifies a method for fully integrating general relativity in fitting the observed spectra from XRBs while upholding the fundamental assumptions and simplicity of the Shakura-Sunyaev model.
@@
Cygnus X-1 is one of the most well-studied and observed XRBs in our Galaxy. Measurements of its spectral distribution in the soft state by RXTE, ASCA, and OSSE in 1996 (Ling et al., 1997; Dotani et al., 1997) and more recently by XMM-Newton, NuSTAR, NICER, and IXPE (Duro, R. et al., 2011; Walton et al., 2016; König, Ole et al., 2024; Steiner et al., 2024) indicate a thermal component from the accretion disk and a hard tail up to the MeV scale. The latter is assumed to originate from the up-scattering of seed photons from the disk in a thermal/non-thermal corona hybrid.
@@
In the remainder of this work, we discuss the disk geometry and the radial temperature obtained from the two PN solutions referring to rotating black holes in Sec. 2. In Sec. 3, we discuss the radiative efficiency perceived by an observer and how it can be employed to acquire an expression for the location of the disk’s inner edge as a function of the spin and the viewing angle. Then, in Sec. 4, we compare our approach to general relativistic results and apply this method to the multi-wavelength spectra detected from Cygnus X-1 in its two primary spectral states. Finally, we conclude our results in Sec. 5.
@@ 2.2 The pseudo-Newtonian solution to the standard accretion disk
where $r=R/R_{g}$, $r_{h}=1+\sqrt{1-\alpha_{*}^{2}}$ is the event horizon in the Kerr metric, and $\beta=r_{in}/r_{h}-1$ denotes a dimensionless constant for a given specific black hole angular momentum. The corresponding potential reproduces pretty accurately the location of the marginally stable orbit and the disk radiative efficiency as predicted by general relativity.
@@ 2.3 Spin-dependent disk temperature profiles
In comparison, Eq. (6) produces negligibly larger disk luminosity across the spectrum for energies below 1 keV that drops slightly more abruptly for higher energies. The deviation is marginal for large $\alpha_{*}$ in the co-rotation case, while they coincide for the rest of the spin values. Both expressions reduce to the temperature function of Gierliński et al. (1999) that implements the Paczyński-Wiita potential.
@@
Extreme Kerr black holes (i.e., $\alpha_{*}>0.9$), governed by frame-dragging effects near their event horizon, are verified as central objects for an increasing number of well-known XRBs (McClintock et al., 2006; Stuchlík & Kološ, 2016; Zhao et al., 2021). Therefore, the spin-dependent profiles given by Eqs. (6) and (9) yield a more realistic measurement of the energy released near those sources.
@@ 3 Spectral angular distribution via disk surface fluctuations
The detectable spectra from accretion disks are heavily dependent on the observer’s viewing angle. Ignoring all relativistic effects, the observed radiative efficiency is $\eta_{rad}^{obs}=2cos\theta\eta_{rad}(R_{ISCO})$, where $\eta_{rad}=L_{disk}/\dot{M}c^{2}$. Regardless, frame-dragging, gravitational redshift, Doppler boost, and light-bending effects significantly alter the angular dependence of the perceived emission flows near the black hole. Hence, the emerging emission pattern following the relativistic kerrbb model is determined by an analytic expression derived by Campitiello et al. (2018) as $\eta_{rad}^{obs}=f(\theta,\alpha_{*})\eta_{rad}(R_{ISCO})$. The function $f(\theta,\alpha_{*})$ is given in the Appendix B.
The corresponding observed total luminosity normalized by $\dot{M}c^{2}$ is plotted for varying values of $\alpha_{*}$ in Fig. 4. We also include the radiative efficiency produced with both pseudo-potentials discussed in this work and the respective theoretical predictions of Novikov-Thorne. The comparison between the PN adaptation and general relativity reveals an increasing deviation for highly rotating black holes.
@@
Figure 4: The solid lines correspond to the radiative efficiency given by the analytic counterpart of the kerrbb model as observed for different system inclinations and black hole spin values (referring to Campitiello et al. (2018) as C18). The purple dashed line showcases the angular dependence corresponding to a Shakura-Sunyaev disk around a static black hole, neglecting all relativistic contributions. The colored dotted lines for each spin value depict the respective theoretical predictions of Novikov & Thorne (1973). For comparison, the respective black and orange-dotted lines represent the Kerr pseudo-Newtonian treatment of this paper.
@@ 3.1 Modification of the disk’s inner boundary
The radiative efficiency is independent of the accretion rate and the spectral hardening factor $f_{col}$. It depends negligibly on the black hole’s stellar mass $M_{bh}$ and the disk’s inner radius $R_{in}$, as long as the outer radius is large enough, $R_{out}>10^{11}\ cm$, akin to the default value set in kerrbb. Fig. 5 illustrates this correlation, which is established by employing Eqs. (6) and (9) in calculating the thermal spectra from the disk. In this work, we fit the power-law form of the Artemova solution (black dotted lines in Fig. 5) with the following expression
@@
The respective orange solid lines of the Mukhopadhyay potential are fitted using a power-law index of $p=0.75$ instead. A noticeable deviation for all spin values, especially near the ISCO radius, deter us from further exploiting the Mukhopadhyay pseudo-force. Also, the fact that the Artemova potential works for radii significantly smaller than $R_{ISCO}$ makes it a very suitable tool for matching the fully relativistic spectrum for a given angle of view.
@@
Furthermore, including all the effects of general relativity and incorporating the correct angular distribution urges the equivalence between Eq. (14) and the observed efficiency given by the analytic function of Campitiello et al. (2018) (see Appendix B). By doing so, we find an expression for the modified inner boundary of the accretion disk, which is given by
@@
This way, equivalent energy distributions to the fully relativistic spectra are reproduced by a standard disk model with the temperature profile of Eq. (6) and an innermost radius given by Eq. (15). The latter is plotted in Fig. 6 and depends on the inclination and angular momentum of the accreting compact object. Basically, when we modify the disk inner radius in order to achieve the kerrbb observed radiative efficiency, we produce spectra that depend on the angle $\theta$ in the same way as dictated by general relativity.
Fig. 6 shows that, independently of the rotation, the physical processes associated with an accretion disk observed from an angle of approximately $\theta=60^{\circ}-65^{\circ}$ identify with the Newtonian-like dynamics of the Shakura-Sunyaev model. Thus, a minimum shift of $R_{in}$ is required to fit the corresponding kerrbb spectra. Consequently, the boundary-modification approach is exceptionally accurate regarding this angle range.
@@
Figure 6: Depending on the source’s viewing angle and the black hole’s rotation, the shifted innermost edge of the accretion disk yields spectra similar to those obtained with the kerrbb model.
4 Results and discussion
4.1 Comparison to General Relativity
There are three primary sources of error concerning the approach presented in this work for calculating relativistic spectra. The first is the intrinsic error margin of the analytic function $f(\theta,\alpha_{*})$, which is relatively small and does not exceed 1% (Campitiello et al., 2018). The second one is due to the PN-potential adoption, which can reach 8-9% in the case of the Artemova solution. Lastly, the error margin introduced in the context of this work is due to the deviation between the PN-generated $\eta_{rad}(R_{in})$ and Eq. (14), with the choice of the $p=1.17$ index, which is also evident in Fig. 5.
@@ 4.1 Comparison to General Relativity
[…]
In the left panel of Fig. 9, we compare the observed disk spectra corresponding to the PN solution of Eq. (6) with the spectra obtained through this work’s boundary-shifting approach according to Eq. (15). We assume a static ($\alpha_{*}=0$), and a rapidly rotating black hole ($\alpha_{*}=0.998$) for greater variety. We can see that the two different approaches are rather close for both spin values when the disk is viewed with an angle of $\theta=60^{\circ}$, a result also verified by Fig. 6 where $R_{in}\approx R_{ISCO}$. However, the significant deviation is obvious when $\theta=0^{\circ}$, especially in the maximum spin case where the disk’s inner edge is closer to the black hole and general relativistic effects are very important.
The PN potential clearly overestimates the total disk luminosity for $\theta\lesssim 60^{\circ}$, while it underestimates it when $\theta\gtrsim 60^{\circ}$. Besides the potential’s intrinsic error, the main reason is that while it offers a very good measurement of the additional energy radiated away by the disk, it cannot offer any information on the spectra a distant observer receives depending on $\theta$. Instead, the PN solution adopts the angular dependence of the classic Shakura-Sunyaev model, $\eta_{rad}^{obs}=2cos\theta\eta_{rad}$.
Our results show an excellent agreement with the fully relativistic formalism of Novikov and Thorne incorporated in the kerrbb model, as seen in the right panel of Fig. 9. The deviation, partially due to the $\eta_{rad}$ error margin of Fig. 8, is minimal even for highly-inclined accretion disks around rapidly rotating central objects.
Furthermore, the approach of this work presents a direct alternative solution in calculating the observed fully relativistic disk spectra without any complex mathematical formalism and the need for a ray-tracing method. Given that the Novikov-Thorne model has been proven very efficient by GRMHD simulations (Kulkarni et al., 2011), our approach could provide a very suitable and straightforward way for estimating black hole spins via the continuum-fitting method.
[…]
Figure 9: Comparison between the disk spectra produced through the boundary-shifting method of this work and (left panel) the Artemova pseudo-Newtonian potential, (right panel) the general relativistic treatment of the kerrbb model. The parameters used are $M_{bh}=10M_{\odot}$, $\dot{M}=10^{18}\ g/s$, and $f_{col}=1$.
@@
References. Miller-Jones et al. (2021), Brocksopp et al. (1999), Krawczynski et al. (2022), Stirling et al. (2001), Gou et al. (2014), Walton et al. (2016), Tetarenko et al. (2019)
4.2 Cygnus X-1: high/soft spectral state
Using Eqs. (6), (15), and the recently redefined Cygnus X-1 parameters of Table 1, we showcase the spin-dependent disk spectra while the system is going through the soft state in Fig. 10. The solid lines correspond to the alignment between the black hole and accretion disk rotational directions. It should be mentioned that those results were obtained for a fixed mass accretion rate of $\dot{M}=8.5\times 10^{-8}$ $M_{\odot}$/yr, which, in most cases, is a free model parameter. In addition, we estimate the radiative efficiency of an extreme rotating black hole to be $\eta_{rad}\approx 0.35$, which agrees with the theoretical predictions. Moreover, employing this work’s approach allows us to constrain more accurately the black hole’s accretion rate as long as we have a precise value of the spin or vice versa.
@@ 4.2 Cygnus X-1: high/soft spectral state
In Fig. 11, we outline the data obtained during observations of the soft spectral state of Cygnus X-1 (Ling et al., 1997; Dotani et al., 1997; Walton et al., 2016). The thermal emission at lower energies is fitted by a black-body spectrum emitted by the inner disk regions. The respective spin is found to be $\alpha_{*}=0.97$, satisfying the lower limit $\alpha_{*}>0.9696$ set by Miller-Jones et al. (2021) for the slightly misaligned spin-orbit case. This value is close to the limits set by the continuum-fitting method as $\alpha_{*}>0.983$ (Gou et al., 2014) and the relativistic reflection modeling as $0.93\gtrsim\alpha_{*}\lesssim 0.96$ (Walton et al., 2016). The high spin yields an accretion disk close to the compact object (i.e., $R_{in}\approx R_{g}$), a primary characteristic of the high/soft state.
@@
$\dot{m}$ 0.007 0.004 mass accretion rate
@@
$T_{cor}$ [keV] 50 150 corona temperature
$\tau_{cor}$ 0.29 0.50 thermal corona optical depth
p 4.9 1.8 corona power-law index
$\gamma_{min}$ 1.8 1.7 minimum energy of non-thermal electrons
@@
h 0.050 0.294 non-thermal/thermal electrons ratio in the corona
@@
$L_{p}/L_{e}$ 0.009 ratio of the energy carried by hadrons to leptons
@@
Moreover, the hard tail above 5 keV is reproduced by the inverse-Compton (IC) scattered disk emission off a thermal/non-thermal corona hybrid described by the parameters of Table 2. In this case, the corona is much more weakly present compared to the hard/low state. The power-law spectrum indicates an electron (or positron) distribution injection, differentiating the two leading emission mechanisms near 10 keV. The injected distribution could be a product of an intense photon annihilation process above the disk level that softens by gradually losing energy due to IC scattering. The resulting power-law has an index of $p>3$ since $p_{inj}>2$. We neglect Compton reflection of the scattered spectrum and irradiation effects beyond those integrated into the kerrbb model as they overcome the purpose of this work. As a result, there is a slight deviation near 6-7 keV in Fig. 11 where the iron $K$ line is detected (Walton et al., 2016).
We assume the injection of 50 non-thermal electrons blending in a thousand thermal ones. Those particles preserve their initial power-law shape, while the rest are described by a Maxwell–Jüttner distribution, which dominates for energies below $\gamma=1.8$. The optical depth in Table 2 corresponds to the thermal counterpart of the corona hybrid. Assuming a fixed spherical region, injecting power-law particles increases corona’s opacity by a factor of $1+h$, where $h$ is the fraction of accelerated to thermal electrons (or positrons) (i.e., $\tau_{cor}^{h}=\tau_{cor}(1+h)\approx 0.3$ in our case).
@@
Figure 11: The high/soft spectral state of Cygnus X-1 was observed by RXTE, ASCA, and NuSTAR (Ling et al., 1997; Dotani et al., 1997; Walton et al., 2016). The observational data are fitted by a black-body spectrum emitted by the inner disk region and an IC scattered component of the disk emission off a combination of thermal and non-thermal corona electrons.